# A Comprehensive Survey for Hyperspectral Image Classification: The Evolution from Conventional to Transformers

### v3

@@ 6.2 Spatial CNN
Pooling layers, like max or average pooling, commonly feature in spatial CNNs to condense feature maps and capture key spatial information. Additional layers, including FC layers, are often integrated to grasp high-level spatial features and enhance classification. In studies by Li et al. [135], Haut et al. [136], and Xu et al. [137], PCA-derived PCs were employed to infuse refined spatial information into CNN frameworks, demonstrating approaches to bolster DL models for HSC. Wang et al. [138] proposed a novel probabilistic neighborhood pooling-based attention network (PNPAN) aiming to enhance model robustness and combat overfitting in HSC. These strategies, including PCA-driven spatial information, random patch networks, and innovative techniques like PNPAN, seek to address challenges linked to limited training samples and overfitting in DL models for HSC.
Ding et al. [139] extracted patches from 2D input images (representing different spectral bands) to train a 2D-CNN architecture. Data-adaptive kernels were learned to capture pertinent features from HS data. Chen et al. [140] proposed a 2D-CNN augmented with Gabor filtering to address overfitting concerns, effectively preserving spatial details. Zhu et al. [141] introduced a deformable HSC network using adaptive deformable sampling locations based on spatial features, enhancing the extraction of complex structures. These methods leverage data-adaptive kernels, handcrafted features like Gabor filters, and deformable sampling locations to improve HSC by capturing relevant spatial information and overcoming challenges such as overfitting with limited training samples.
Advantages: Spatial CNN captures vital spatial-contextual information for HSC, facilitating the understanding of spatial patterns and relationships. This capability is crucial for handling spatially varying features in HSIs, including structures, textures, and localized objects. Another strength lies in their ability to fuse spatial and spectral features, enhancing classification accuracy by leveraging the complementary aspects of these dimensions.
@@ 6.3 Spectral-Spatial CNN
Integrating spatial features with spectral information enhances spectral-spatial pixel-wise HSC. For instance, Ran et al. [142] introduced spatial pixel pair features (PPF), emphasizing immediate neighborhood pairs with identical labels for enhanced discriminative power. Zhong et al. [143] proposed a Spectral-Spatial Residual Network (SSRN) utilizing 3D-convolutions in spectral and spatial residual blocks to extract joint representations. Paoletti et al. [144] presented a 3D CNN framework effectively capturing spectral-spatial relationships for improved classification accuracy. Li et al. [145] employed adaptive weight learning for dynamic adjustment of spatial information importance, capturing variations in spatial context. Roy et al. [146] proposed adaptively determined adjustable receptive fields, enhancing joint feature extraction in a spectral-spatial residual network. Additionally, Roy et al. [147] fused maximum and minimum convolutional features, reported to enhance classification performance in HSC tasks.
@@
Paoletti et al. [148] addressed 3D CNN limitations in exploiting rotation equivariance by introducing translation-equivariant representations, enhancing spatial feature robustness in HSC tasks. Zhang et al. [149] proposed an end-to-end 3D lightweight CNN to counter overfitting and gradient vanishing challenges with limited labeled samples. Jia et al. [150] introduced spatial-spectral Schroedinger eigenmaps (SSSE) to balance numerous trainable parameters and scarce labeled samples. Roy et al. [151] incorporated a lightweight bag-of-feature paradigm into a spectral-spatial squeeze-and-excitation residual network. Roy et al. [152] introduced Morphological CNN (MorphCNN), combining spectral and spatial features for effective information utilization.
@@
Li et al. [145] proposed a two-stage framework with adaptive learning of input patch weights in the first stage, extracting joint shallow features and obtaining deep hierarchical features through a Stacked AE (SAE) network. Classification is performed using a Multinomial Logistic Regression (MLR) layer. Li et al. [153] introduced a 3D-CNN model, comparing its performance with spectral-based Deep Belief Networks (DBN), SAE, and 2D-spatial CNNs. Roy et al. [154] presented a bilinear fusion mechanism combining global pooling and max-pooling branches. Jiao et al. [155] introduced a deep multiscale spectral-spatial feature extraction approach, learning discriminative features from spatially diverse images using a Fully Convolutional Network (FCN). These spatial features are fused with spectral information using a weighted fusion strategy for accurate pixel-wise classification.
@@
He et al. [156] used PCA-transformed images to generate multi-scale cubes, extracting handcrafted features with multi-scale covariance maps to capture spectral-spatial details. Zhang et al. [157] implemented a dual-channel CNN framework, employing a 1D-CNN for spectral and a 2D-CNN for spatial hierarchical feature extraction, combining both for classification. He et al. [158] introduced a multiscale 3D deep CNN for end-to-end HSC, jointly learning 1D spectral and 2D multiscale spatial features without pre-processing like PCA. Dong et al. [159] embedded a band attention module within a CNN to selectively focus on informative bands, reducing redundancy and noise in HSC. Cheng et al. [160] combined CNN with a metric learning-based framework, using CNN for deep spatial information extraction and metric learning for spectral and spatial feature fusion. Gong et al. [161] merged a multi-scale convolution-based CNN with diversified deep metrics based on determinantal point process (DPP) priors [162], employing multi-scale filters and DPP-based diversified metrics for enhanced HSI representational ability.
@@
Liu et al. [163] introduced an HSC framework for multi-scale spatial feature extraction, constructing a three-channel RGB image from HSIs to leverage existing networks. Sellami et al. [164] presented an adaptive band selection-based semi-supervised 3D CNN, jointly exploiting spectral-spatial features. Ma et al. [165] proposed a two-branch Deep CNN, with one branch extracting spatial information and the other utilizing a contextual DNN for spectral features. Mei et al. [166] used a 3D convolutional AE for unsupervised simultaneous exploitation of spectral-spatial features. Roy et al. [167] explored a dual-attention-based AE-decoder network for unsupervised HS band selection and joint feature extraction. Roy et al. [168] introduced a lightweight HetConv3D for HSC with noisy labels, combining spectral and spatial kernel features. Paoletti et al. [169] presented a separable attention network dividing input feature maps for global contextual information encoding. Roy et al. [170] introduced G2C-Conv3D to combine intensity-level semantic information and gradient-level detailed information during convolution operations.
@@
6.4 Graph Convolutional Networks for HSC
Graph Convolutional Networks (GCNs) have garnered attention for their adeptness in processing non-grid high-dimensional data and their adaptable network architecture [171]. These characteristics open up possibilities for effective handling of HS data [172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182]. GCNs are particularly proficient in capturing relationships between data or samples, making them well-suited for modeling spatial relationships among spectral signatures in HSIs, as depicted in Figure 8. However, the construction of large graphs poses challenges due to computational demands, limiting the popularity of GCNs in HSC compared to CNNs. Despite this, exploratory studies have begun to leverage GCNs in certain HSC tasks [183].
@@ 6.4 Graph Convolutional Networks for HSC
By incorporating graph structures and spatial relationships, GCNs provide a robust framework for HSC, capturing intricate interactions between spectral signatures and leveraging spatial context for improved accuracy. Despite their potential, it is crucial to address computational complexities and scalability issues with large HS datasets. Researchers have proposed strategies to mitigate these challenges. For instance, a second-order GCN in [184] models spatial-spectral relations on manifolds [185] to reduce computational costs, while [186] utilizes superpixel segmentation to enhance GCN efficiency in handling larger pixel sets for land cover classification. However, fundamental issues persist. Hong et al. proposed miniGCN in [183], training GCNs in a mini-batch fashion akin to CNNs, effectively reducing computational costs and enabling quantitative comparisons and fusion with CNNs, leading to the development of FuNet for HSC.
@@
6.5 Future Research Directions
@@ 7 Deep Belief Network for HSC
Introduced by Hinton et al. in 2006 [187], the Deep Belief Network (DBN) is a hierarchical DNN that utilizes unsupervised learning to sequentially learn features from input data as shown in Figure 9. Constructed with multiple layers using Restricted Boltzmann Machines (RBMs), RBMs are two-layer neural networks where visible units connect to hidden units [188]. RBMs play a crucial role in extracting informative features from input data, forming a layer-by-layer architecture in DBNs. This hierarchical structure enables greedy training, with each RBM capturing progressively abstract features from HS data.
@@
Widely applied in HSC, DBNs, such as in Ayhan et al.’s work [189], have been utilized for land cover classification, emphasizing the integration of spectral and spatial information. The typical training process involves unsupervised pre-training with unlabeled samples, followed by supervised fine-tuning using labeled samples. However, this approach may lead to co-adaptation issues among hidden units and activation neuron sparsity, limiting the diversity of learned features. Addressing these challenges [190, 191], Zhong et al. proposed a diversified DBN model [192]. This model introduces a regularization technique during both pre-training and fine-tuning stages, promoting diversity among hidden units and activation neurons. By mitigating co-adaptation and enhancing activation selectivity, the diversified DBN model provides a potential solution to improve DBN classification performance in HSC, enhancing representation power and discriminative capabilities.
In HSC, researchers have introduced diverse strategies for efficient texture feature extraction. Li et al. proposed a DBN-based texture feature enhancement framework [193], incorporating band grouping, sample band selection, and guided filter techniques to improve texture features. Another approach by Tan et al. implemented a parallel layers framework [194], using a Gaussian-Bernoulli RBM for high-level, local invariant, and nonlinear feature extraction, followed by a logistic regression layer for classification. Both methods utilize DBN-based frameworks and parallel layers architectures to enhance texture features, emphasizing the importance of effective feature extraction for accurate HSC.
To improve classification accuracy, studies have focused on integrating spectral and spatial information [193]. Li et al. presented a DBN framework with a logistic regression layer, showcasing that joint exploitation of spectral-spatial features enhances accuracy [195]. Similarly, Sellami et al. proposed a spectral-spatial graph-based RBM method [196], constructing a spectral-spatial graph to measure similarity based on both spectral and spatial details. The RBM extracted joint spectral-spatial features, contributing to improved classification performance when integrated into a DBN and logistic regression layer. These approaches highlight the significance of considering both spectral and spatial information for effective HSC.
@@ 8 Recurrent Neural Networks for HSC
RNNs, featuring loop connections that link current and previous steps, excel in learning from temporal sequences [197]. In the context of HS data, RNNs treat spectral bands as sequential time steps, providing a temporal processing advantage [198]. Three fundamental RNN models include Vanilla, LSTM, and GRU (Figure 10), each offering distinct strategies to address challenges related to learning and capturing long-term dependencies in sequential data. While Vanilla RNNs are straightforward, they may struggle with extended sequences, prompting the use of LSTM and GRU models designed to mitigate the vanishing gradient problem and handle long-term dependencies more effectively. Leveraging the temporal aspects of RNNs for sequential processing in HSI data allows researchers to harness their power in capturing and modeling temporal dynamics.
Figure 10: Internal architecture of LSTM and GRU [1]
@@
Hang et al. [199] proposed an RNN framework using the novel parametric rectified tanh activation function and GRU for sequential HS data analysis. Zhang et al. [75] introduced an RNN-based local spatial sequential (LSS) method, fusing Gabor filters and differential morphological profiles for low-level feature extraction, followed by RNN modeling for high-level feature extraction. Zhou et al. [200] employed separate LSTM networks for spectral and spatial features, combining their outputs using a decision fusion strategy. Sharma et al. [201] utilized LSTM cells for multi-temporal and multi-spectral information, capturing complex patterns and dependencies in HSI for effective land cover classification.
Wu et al. [130] introduced a convolutional RNN model for HSC, using initial convolutional layers to extract position-invariant middle-level features and recurrent layers to capture spectral-contextual details. In a semi-supervised setting, Wu et al. [202] applied a similar CRNN model, incorporating pseudo-labels for leveraging unlabeled data. Zhou et al. [203] proposed an integrated CNN and GRU-based fusion network, where CNN extracts spatial features and GRU performs both feature-level and decision-level fusion. Luo et al. [204] introduced a CNN-parallel GRU-based RNN architecture, simplifying GRU training and enhancing performance by combining it with the CNN component. Liu et al. [205] presented the Bidirectional CLSTM model, jointly exploiting spectral-spatial features of HSI. Shi et al. [206] combined multiscale local spectral-spatial features from a 3D CNN with a hierarchical RNN. Yang et al. [207] proposed recurrent 2D CNN and 3D CNN models for HSC. Hang et al. [199] introduced a cascade RNN architecture with two layers of GRU-based RNNs, reducing redundant spectral bands in the first layer and learning features in the second layer.

### v4

@@ 6 Convolutional Neural Networks for HSC
Pooling layers, like max or average pooling, commonly feature in spatial CNNs to condense feature maps and capture key spatial information. Additional layers, including FC layers, are often integrated to grasp high-level spatial features and enhance classification. In studies by Li et al. (Li et al., 2018b, ), Haut et al. (Haut et al.,, 2019), and Xu et al. Xu et al., (2018), PCA-derived PCs were employed to infuse refined spatial information into CNN frameworks, demonstrating approaches to bolster DL models for HSC. Wang et al. (Wang et al.,, 2021) proposed a novel probabilistic neighborhood pooling-based attention network (PNPAN) aiming to enhance model robustness and combat overfitting in HSC. These strategies, including PCA-driven spatial information, random patch networks, and innovative techniques like PNPAN, seek to address challenges linked to limited training samples and overfitting in DL models for HSC.
Ding et al. (Ding et al.,, 2017) extracted patches from 2D input images (representing different spectral bands) to train a 2D-CNN architecture. Data-adaptive kernels were learned to capture pertinent features from HS data. Chen et al. (Chen et al.,, 2017) proposed a 2D-CNN augmented with Gabor filtering to address overfitting concerns, effectively preserving spatial details. Zhu et al. (Zhu et al., 2018a, ) introduced a deformable HSC network using adaptive deformable sampling locations based on spatial features, enhancing the extraction of complex structures. These methods leverage data-adaptive kernels, handcrafted features like Gabor filters, and deformable sampling locations to improve HSC by capturing relevant spatial information and overcoming challenges such as overfitting with limited training samples.
Advantages: Spatial CNN captures vital spatial contextual information for HSC, facilitating the understanding of spatial patterns and relationships. This capability is crucial for handling spatially varying features in HSIs, including structures, textures, and localized objects. Another strength lies in their ability to fuse spatial and spectral features, enhancing classification accuracy by leveraging the complementary aspects of these dimensions.
@@
Integrating spatial features with spectral information enhances spectral-spatial pixel-wise HSC. For instance, Ran et al. (Ran et al.,, 2017) introduced spatial pixel pair features (PPF), emphasizing immediate neighborhood pairs with identical labels for enhanced discriminative power. Zhong et al. (Zhong et al., 2018b, ) proposed a Spectral-Spatial Residual Network (SSRN) utilizing 3D-convolutions in spectral and spatial residual blocks to extract joint representations. Paoletti et al. (Paoletti et al., 2018a, ) presented a 3D CNN framework effectively capturing spectral-spatial relationships for improved classification accuracy. Li et al. (Li et al., 2019b, ) employed adaptive weight learning for dynamic adjustment of spatial information importance, capturing variations in spatial context. Roy et al. (Roy et al., 2021d, ) proposed adaptively determined adjustable receptive fields, enhancing joint feature extraction in a spectral-spatial residual network. Additionally, Roy et al. (Roy et al., 2021f, ) fused maximum and minimum convolutional features, reported to enhance classification performance in HSC tasks.
@@
Paoletti et al. (Paoletti et al., 2020b, ) addressed 3D CNN limitations in exploiting rotation equivariance by introducing translation-equivariant representations, enhancing spatial feature robustness in HSC tasks. Zhang et al. (Zhang et al., 2019a, ) proposed an end-to-end 3D lightweight CNN to counter overfitting and gradient vanishing challenges with limited labeled samples. Jia et al. (Jia et al.,, 2020) introduced spatial-spectral Schroedinger eigenmaps (SSSE) to balance numerous trainable parameters and scarce labeled samples. Roy et al. (Roy et al., 2020a, ) incorporated a lightweight bag-of-feature paradigm into a spectral-spatial squeeze-and-excitation residual network. Roy et al. (Roy et al., 2021e, ) introduced Morphological CNN (MorphCNN), combining spectral and spatial features for effective information utilization.
@@
Li et al. (Li et al., 2019b, ) proposed a two-stage framework with adaptive learning of input patch weights in the first stage, extracting joint shallow features and obtaining deep hierarchical features through a Stacked AE (SAE) network. Classification is performed using a Multinomial Logistic Regression (MLR) layer. Li et al. (Li et al., 2017b, ) introduced a 3D-CNN model, comparing its performance with spectral-based Deep Belief Networks (DBN), SAE, and 2D-spatial CNNs. Roy et al. (Roy et al., 2020c, ) presented a bilinear fusion mechanism combining global pooling and max-pooling branches. Jiao et al. (Jiao et al.,, 2017) introduced a deep multiscale spectral-spatial feature extraction approach, learning discriminative features from spatially diverse images using a Fully Convolutional Network (FCN). These spatial features are fused with spectral information using a weighted fusion strategy for accurate pixel-wise classification.
@@
He et al. (He et al.,, 2018) used PCA-transformed images to generate multi-scale cubes, extracting handcrafted features with multi-scale covariance maps to capture spectral-spatial details. Zhang et al. (Zhang et al., 2017a, ) implemented a dual-channel CNN framework, employing a 1D-CNN for spectral and a 2D-CNN for spatial hierarchical feature extraction, combining both for classification. He et al. (He et al., 2017a, ) introduced a multiscale 3D deep CNN for end-to-end HSC, jointly learning 1D spectral and 2D multiscale spatial features without pre-processing like PCA. Dong et al. (Dong et al.,, 2019) embedded a band attention module within a CNN to selectively focus on informative bands, reducing redundancy and noise in HSC. Cheng et al. (Cheng et al.,, 2018) combined CNN with a metric learning-based framework, using CNN for deep spatial information extraction and metric learning for spectral and spatial feature fusion. Gong et al. (Gong et al.,, 2019) merged a multi-scale convolution-based CNN with diversified deep metrics based on determinantal point process (DPP) priors (Zhong et al.,, 2015), employing multi-scale filters and DPP-based diversified metrics for enhanced HSI representational ability.
@@
Liu et al. (Liu et al., 2020a, ) introduced an HSC framework for multi-scale spatial feature extraction, constructing a three-channel RGB image from HSIs to leverage existing networks. Sellami et al. (Sellami et al.,, 2019) presented an adaptive band selection-based semi-supervised 3D CNN, jointly exploiting spectral-spatial features. Ma et al. (Ma et al.,, 2018) proposed a two-branch Deep CNN, with one branch extracting spatial information and the other utilizing a contextual DNN for spectral features. Roy et al. (Roy et al., 2020b, ) explored a dual-attention-based AE-decoder network for unsupervised HS band selection and joint feature extraction. Roy et al. (Roy et al., 2021b, ) introduced a lightweight HetConv3D for HSC with noisy labels, combining spectral and spatial kernel features. Paoletti et al. (Paoletti et al.,, 2021) presented a separable attention network dividing input feature maps for global contextual information encoding. Roy et al. (Roy et al., 2021c, ) introduced G2C-Conv3D to combine intensity-level semantic information and gradient-level detailed information during convolution operations.
@@
6.d Graph Convolutional Networks for HSC
Graph Convolutional Networks (GCNs) have garnered attention for their adeptness in processing non-grid high-dimensional data and their adaptable network architecture (Kipf and Welling,, 2016). These characteristics open up possibilities for effective handling of HS data (Yu et al.,, 2023; Ding et al.,, 2024). GCNs are particularly proficient in capturing relationships between data or samples, making them well-suited for modeling spatial relationships among spectral signatures in HSIs, as depicted in Figure 8. However, the construction of large graphs poses challenges due to computational demands, limiting the popularity of GCNs in HSC compared to CNNs. Despite this, exploratory studies have begun to leverage GCNs in certain HSC tasks (Hong et al., 2021a, ).
@@
By incorporating graph structures and spatial relationships, GCNs provide a robust framework for HSC, capturing intricate interactions between spectral signatures and leveraging spatial context for improved accuracy. Despite their potential, it is crucial to address computational complexities and scalability issues with large HS datasets. Researchers have proposed strategies to mitigate these challenges. For instance, a second-order GCN in (Qin et al.,, 2018) models spatial-spectral relations on manifolds (Hong et al., 2019c, ) to reduce computational costs, while (Wan et al.,, 2019) utilizes superpixel segmentation to enhance GCN efficiency in handling larger pixel sets for land cover classification. However, fundamental issues persist. Hong et al. proposed miniGCN in (Hong et al., 2021a, ), training GCNs in a mini-batch fashion akin to CNNs, effectively reducing computational costs and enabling quantitative comparisons and fusion with CNNs, leading to the development of FuNet for HSC.
@@
6.e Future Research Directions
@@ 7 Deep Belief Network for HSC
Introduced by Hinton et al. in 2006 (Hinton et al.,, 2006), the Deep Belief Network (DBN) is a hierarchical DNN that utilizes unsupervised learning to sequentially learn features from input data as shown in Figure 9. Constructed with multiple layers using Restricted Boltzmann Machines (RBMs), RBMs are two-layer neural networks where visible units connect to hidden units (Zhang et al., 2018a, ). RBMs play a crucial role in extracting informative features from input data, forming a layer-by-layer architecture in DBNs. This hierarchical structure enables greedy training, with each RBM capturing progressively abstract features from HS data.
@@
Widely applied in HSC, DBNs, such as in Ayhan et al.’s work (Ayhan and Kwan,, 2017), have been utilized for land cover classification, emphasizing the integration of spectral and spatial information. The typical training process involves unsupervised pre-training with unlabeled samples, followed by supervised fine-tuning using labeled samples. However, this approach may lead to co-adaptation issues among hidden units and activation neuron sparsity, limiting the diversity of learned features. Addressing these challenges (Shaham et al.,, 2016; Xiong et al.,, 2015), Zhong et al. proposed a diversified DBN model (Zhong et al.,, 2017). This model introduces a regularization technique during both pre-training and fine-tuning stages, promoting diversity among hidden units and activation neurons. By mitigating co-adaptation and enhancing activation selectivity, the diversified DBN model provides a potential solution to improve DBN classification performance in HSC, enhancing representation power and discriminative capabilities.
In HSC, researchers have introduced diverse strategies for efficient texture feature extraction. Li et al. proposed a DBN-based texture feature enhancement framework (Li et al., 2018a, ), incorporating band grouping, sample band selection, and guided filter techniques to improve texture features. Both methods utilize DBN-based frameworks and parallel layers architectures to enhance texture features, emphasizing the importance of effective feature extraction for accurate HSC.
To improve classification accuracy, studies have focused on integrating spectral and spatial information (Li et al., 2018a, ). Li et al. presented a DBN framework with a logistic regression layer, showcasing that joint exploitation of spectral-spatial features enhances accuracy (Li et al., 2019a, ). Similarly, Sellami et al. proposed a spectral-spatial graph-based RBM method (Sellami and Farah,, 2019), constructing a spectral-spatial graph to measure similarity based on both spectral and spatial details. The RBM extracted joint spectral-spatial features, contributing to improved classification performance when integrated into a DBN and logistic regression layer. These approaches highlight the significance of considering both spectral and spatial information for effective HSC.
@@ 8 Recurrent Neural Networks for HSC
RNNs, featuring loop connections that link current and previous steps, excel in learning from temporal sequences (Williams and Zipser,, 1989). In the context of HS data, RNNs treat spectral bands as sequential time steps, providing a temporal processing advantage (Paoletti et al., 2020a, ). Three fundamental RNN models include Vanilla, LSTM, and GRU (Figure 10), each offering distinct strategies to address challenges related to learning and capturing long-term dependencies in sequential data. While Vanilla RNNs are straightforward, they may struggle with extended sequences, prompting the use of LSTM and GRU models designed to mitigate the vanishing gradient problem and handle long-term dependencies more effectively. Leveraging the temporal aspects of RNNs for sequential processing in HSI data allows researchers to harness their power in capturing and modeling temporal dynamics.
Figure 10: Internal architecture of LSTM and GRU Ahmad et al., 2021c
@@
Hang et al. (Hang et al.,, 2019) proposed an RNN framework using the novel parametric rectified tanh activation function and GRU for sequential HS data analysis. Zhang et al. (Zhang et al., 2018b, ) introduced an RNN-based local spatial sequential (LSS) method, fusing Gabor filters and differential morphological profiles for low-level feature extraction, followed by RNN modeling for high-level feature extraction. Sharma et al. (Sharma et al.,, 2018) utilized LSTM cells for multi-temporal and multi-spectral information, capturing complex patterns and dependencies in HSI for effective land cover classification.
Wu et al. (Wu and Prasad, 2017a, ) introduced a convolutional RNN model for HSC, using initial convolutional layers to extract position-invariant middle-level features and recurrent layers to capture spectral-contextual details. In a semi-supervised setting, Wu et al. (Wu and Prasad, 2017b, ) applied a similar CRNN model, incorporating pseudo labels for leveraging unlabeled data. Luo et al. (Luo,, 2018) introduced a CNN-parallel GRU-based RNN architecture, simplifying GRU training and enhancing performance by combining it with the CNN component. Liu et al. (Liu et al., 2017b, ) presented the Bidirectional CLSTM model, jointly exploiting spectral-spatial features of HSI. Shi et al. (Shi and Pun,, 2018) combined multiscale local spectral-spatial features from a 3D CNN with a hierarchical RNN. Yang et al. (Yang et al.,, 2018) proposed recurrent 2D CNN and 3D CNN models for HSC. Hang et al. (Hang et al.,, 2019) introduced a cascade RNN architecture with two layers of GRU-based RNNs, reducing redundant spectral bands in the first layer and learning features in the second layer.